### v1

@@ V Conclusion
We proposed a speech quality estimator based on unconditional score-based diffusion models trained on clean speech only. Using the natural likelihood computation abilities of score-based models, the proposed estimator can estimate the quality of speech utterances without having any access to paired data. The resulting measure is non-intrusive and yet correlates well with intrusive metrics on noisy speech benchmarks. When evaluating utterances processed by speech enhancement baselines, our method is the only unsupervised non-intrusive metric that properly correlates with other intrusive metrics, as well as supervised DNN-based non-intrusive models.
VI Acknowledgement
We would like to thank J. Berger and Rohde&Schwarz SwissQual AG for their support with POLQA.

### v2

@@ 5 Conclusion
We proposed a speech quality estimator based on unconditional score-based diffusion models trained on clean speech only. Using the natural likelihood computation abilities of score-based models, the proposed estimator can estimate the quality of speech utterances without having any access to paired data. The resulting measure is non-intrusive and yet correlates well with intrusive metrics on noisy speech benchmarks. When evaluating utterances processed by speech enhancement baselines, our method showed the highest correlation with the scores assigned by human listeners.
6 Acknowledgement
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 498394658. We acknowledge the support by the DFG in the transregio project Crossmodal Learning (TRR 169) and DASHH (Data Science in Hamburg – Helmholtz Graduate School for the Structure of Matter) with Grant-No. HIDSS-0002. We would like to thank J. Berger and Rohde&Schwarz SwissQual AG for their support with POLQA.